# Continual Lie algebras determined by chain complexes

### v5

@@ Abstract.
Continual Lie algebras are infinite-dimensional generalizations of Lie algebras with discrete root system by considering continual root systems. In this paper we establish a general relation between chain complexes and continual Lie algebras. The natural orthogonality condition with respect to a product among elements of a chain complex $\mathcal{C}$ spaces brings about to $\mathcal{C}$ the structure of a graded algebra with differential relations. We prove the main result of this paper: a chain complex endowed with an appropriate Leibniz-property product of elements of its spaces brings about the structure of a continual Lie algebra with the root space determined by parameters for the complex. That provides a new source of examples of continual Lie algebras. Finally, as an example, we consider the case of Čech-de Rham complex associated to a foliation of a smooth manifold. In a particular case of this chain complex, we derive explicitly the commutation relations for the corresponding continual Lie algebra.
@@ 1. Introduction
Continual Lie algebras introduced in [16] are generalizations of infinite-dimensional Lie algebras with systems of discrete roots [12]. Then the notion was generalized and developed in [17][20]. These algebras are formulated in such a way that the space of roots is defined by continual sets of vectors. In commutation relations, the generators depend on kernels which are functionals of continual roots. the Jacobi identity for continual Lie algebras result in non-trivial functional relations for kernels. Similar constructions appeared recently in the study of certain Hall algebras of coherent sheaves described in terms of certain continuum limits of Kac-Moody algebras [1].
Though a general theory of continual Lie algebras and their representations is missing, there exist many applications, especially in the theory of completely integrable and exactly solvable models [14]. Various applications can be also found in other fields of mathematics [4, 6, 20].
In original papers [16][20], in order to find new classes of examples of continual Lie algebras various approaches were studied. In a series of papers [18, 19] the authors have generalized first known examples and have found new non-trivial and fundamental ones arising from various branches of mathematics, in particular, functional analysis [20], differential geometry, algebraic geometry, non-commutative geometry, and mathematical physics [4, 8]. Some of them came from physical applications, e.g., equations related to the Ricci flow and cosmology [4, 8]. Though they appear in various corners of modern mathematical research, at some stage, the arsenal of ideas of generation of new examples of continual Lie algebras was exhausted. An intriguing question is how to construct a general way which would allow to create continual Lie algebras. In many cases, this problem reduces to the question of finding new appropriate solutions for relations (1.1) (which is a very interesting problem by itself) following from the Jacobi identity.
The idea of this work is to use properties of general chain complexes, (with complex spaces depending on sets of parameters) in order to derive the structure of continual Lie algebras. The main result of this paper consists in the following
@@
For the set $I_{0}$ of all pairs of independent elements $\alpha_{i}^{(K_{i})}$, the orthogonality condition (2.5), the generators $\left\{\chi,\;\delta(n_{0})\chi,\;\Phi,\;\delta(n)\Phi,\;\alpha_{i}^{(K_{i})},\;\delta(n(i))\alpha_{i}^{(K_{i})}\right\}$, $n(i)=n_{i}^{(K_{i})}$, the relations (2.2), and the Jacobi identity (1.1) form a continual Lie algebra $\mathcal{G}(\Theta_{n(i)})$ with the root space depending on the set of parameters $\Theta_{n(i)}$, of spaces $C\left(n(i),\Theta_{n(i)}\right)$.
@@
The plan of the paper is the following. In the next subsection we recall the notion of a continual Lie algebra [16]. In Section 2, starting from a general (infinite) chain complex, we first define a product acting on elements of various spaces of a complex. We require the Leibniz rule to be fulfilled for (co)boundary operators with respect to this product. In addition to that we assume its skew-symmetry properties. Thus we endow the complex with the structure of a graded differential algebra $\mathcal{C}$.
@@
We then apply the natural condition of orthogonality with respect to a product among the chain complex spaces. For instance, in differential geometry, the orthogonality condition reduces to the integrability condition for differential forms. Further actions of (co)boundary operators, as well as consequences from the definition of the multiplication, generate a system of differential relations for elements of various chain complex spaces with extra compatibility conditions for indices. This determines the structure of a graded algebra with differential relations for $\mathcal{C}$. By choosing independent elements satisfying the above relations, we then show the main result of this paper: for a graded differential algebra associated to a chain complex, the independent generators, the system of differential relations, and the Jacobi identity define the structure of a continual Lie algebra with the root space provided by the spaces parameters of a chain complex.
@@
In Section 3 we consider the example of the Čech-de Rham cochain bicomplex [5] (c.f. subsection 3.1) associated to a smooth manifold foliation [7, 2, 3]. This bicomplex has a deep geometric meaning [9] and is defined for the spaces of differential forms and holonomy embeddings acting between sections of the transversal basis for a foliation. In subsection 3.1 we recall the notions of holonomy mappings, transversal sections, and transversal basis for a foliation. The space of differential forms has already a structure of bigraded differential algebra with respect to the natural product [5]. For the bicomplex under consideration, we define the second product, satisfying properties required for the construction of a system of differential relations. According to the exposition we give in Section 2, by involving the orthogonality conditions for an arbitrary pairs of the Čech-de Rham bicomplex, we apply the general scheme of construction of a bigraded differential algebra. Then we single out generators and commutation relations for the corresponding continual Lie algebra with the space of roots and kernels given by the sets of holonomy embeddings. In a particular geometric case associated to a codimension one foliation over a three-dimensional smooth manifold, we start from the integrability condition applied to elements of the same bicomplex space which leads to the appearance of a continual Lie algebra with generators, kernels, and commutation relations explicitly described.
@@ 1.1. Continual Lie algebras
In this subsection we recall the notion of a continual Lie algebra introduced in [17]. It was then studied in [18, 19]. Suppose $E$ is an associative algebra over $\mathbb{R}$ or $\mathbb{C}$, and $K_{0}$, $K_{\pm 1}$, $K_{0,0}:E\times E\to E$, are bilinear mappings. The local part of a continual Lie algebra can be defined as ${\widehat{\mathcal{G}}}=\mathcal{G}_{-1}\oplus\mathcal{G}_{0}\oplus\mathcal{G}_{+1}$, where $\mathcal{G}_{i}$, $i=0,\pm 1$. The subspaces $\mathcal{G}_{i}$ consist of the elements $\left\{X_{i}(\phi),\phi\in E\right\}$, $i=0,\pm 1$. The generators $X_{i}(\phi)$ are subject to commutation relations. For instance, they can have the form
@@
for all $\phi$, $\psi$, $\chi\in E$. An infinite dimensional algebra $\mathcal{G}\left(E;\widehat{K}\right)=\mathcal{G}^{\prime}\left(E;\widehat{K}\right)/J$, is called a continual contragredient Lie algebra, where $\mathcal{G}^{\prime}(E;\widehat{K})$ is a Lie algebra freely generated by ${\widehat{\mathcal{G}}}$, and $J$ is the largest homogeneous ideal with trivial intersection with $\mathcal{G}_{0}$ (consideration of the quotient is equivalent to imposing the Serre relations in the case of finite-dimensional simple complex Lie algebras) [18, 19]. $\mathcal{G}$ is endowed with a $\mathbb{Z}$-grading $\mathcal{G}=\bigoplus_{n\in\mathbb{Z}}\mathcal{G}_{n}$, where elements of subspaces $\mathcal{G}_{n}$ satisfy the standard grading condition $\left[\mathcal{G}_{n},\mathcal{G}_{m}\right]\subset\mathcal{G}_{n+m}$, where $X_{n}(\phi)\in\mathcal{G}_{n}$, and higher mapping $K_{n,m}(\phi,\psi)$ are present in commutation relations among $X_{n}(\phi)$ and $X_{n}(\psi)$. In general, the Jacobi identity for generators that belong to all grading spaces has the form
@@ 2.1. The graded differential algebra
In this section we formulate the construction which allows us to generate an algebra of graded differential relations starting from a chain complex. Consider a complex of spaces with elements depending on sets of parameters $\Theta_{i}$, $i\in\mathbb{Z}$, and given by
$\ldots\stackrel{{\scriptstyle\delta(i-1)}}{{\longrightarrow}}C(i,\Theta_{i})\stackrel{{\scriptstyle\delta(i)}}{{\longrightarrow}}C(i+1,\Theta_{i+1})\stackrel{{\scriptstyle\delta(i+1)}}{{\longrightarrow}}\ldots,$ (2.1)
@@
so that for any elements $\Phi_{i}\in C(i,\Theta_{i})$ and $\Phi_{j}\in C(j,\Theta_{j})$, $\Phi_{i}\cdot\Phi_{j}=\Phi_{f(i,j)}$, where $f(i,j)$ and $g(i,j)$ are some functions of indices $i$ and $j$. Note that a function $f(i,j)$ defined the index for the product (2.2) resulting space. The function $g(i,j)$ defines the index for the resulting set of parameters. Let us assume that for all $i\in\mathbb{Z}$, the Leibniz rule formula for the operator $\delta(i)$ takes place, i.e.,
$\delta(k+l)\left(\Phi_{k}\cdot\Phi_{l}\right)=\left(\delta(k)\Phi_{k}\right)\cdot\Phi_{l}+(-1)^{\deg\left(\Phi_{k}\right)}\Phi_{k}\cdot\delta(l)\Phi_{l}.$ (2.3)
We assume also that the chain complex above has some properties of an algebra with respect to the product (2.2). Namely, consider the elements $\Phi_{k}\in C(k,\Theta_{k})$, and $\Phi_{l}\in C(l,\Theta_{l})$. Let us assume, in particular, that the product (2.2) satisfies the condition
$\Phi_{k}\cdot\Phi_{l}=-\Phi_{l}\cdot\Phi_{k}.$ (2.4)
for all $k$, $l\in\mathbb{Z}$, $\Phi_{k}\in C(k,\Theta_{k})$ and $\Phi_{l}\in C(l,\Theta_{l})$. Altogether, the complex (2.1) form a graded differential algebra $\mathcal{C}$. In what follows, we skip the sets of parameters $\Theta_{i}$ in the notations for the spaces of complexes, i.e., we set $C(i)=C(i,\Theta_{i})$. Though one has to keep in mind that a product of elements of two spaces of complex (2.1) depends on the resulting set of parameters $\Theta_{g(i,j)}$.
@@ 2.2. The algebra of graded differential relations
Let us assume that the product (2.2) is defined in such a way that common parameters of the sets $\Theta_{i}$ and $\Theta_{j}$ are present in the resulting set $\Theta_{g(i,j)}$ only once. For some $i$ and $j\in\mathbb{Z}$, let us introduce the orthogonality condition for a pair $C(i,\Theta_{i})$ and $C(j,\Theta_{j})$, with respect to the product (2.2). In particular, let us require that for a pair $C(i,\Theta_{i})$, $C(j,\Theta_{j})$, there exist subspaces $C^{\prime}(i)\subset C(i,\Theta_{i})$ and $C^{\prime}(j)\subset C(j,\Theta_{j})$, such that, for any $\Phi_{i}\in C^{\prime}(i)$ and $\Phi_{j}\in C^{\prime}(j)$,
@@
namely, $\Phi_{i}$ is supposed to be orthogonal to $\delta(j)\Phi_{j}$ with respect to (2.2).
By applying further differentials to (2.5) (and further consequences of such action), and using properties of a particular function $f(i,j)$ we obtain relations among elements of spaces $C(i,\Theta_{i})$, $i\in\mathbb{Z}$. In particular, taking into account that both sides of such relations belong to the same space, we obtain limitations (depending on the function $f$) on indices. In differential geometry, the orthogonality condition (2.5) provides the definition of integrability conditions for differential forms, and leads to the Frobenius theorem (see, e.g., [10]).
Let us explain the notations we will use on few next pages. In (2.2) we obtain (infinite) sequences of pairs of differential relations of the form (2.7)–(2.9). The (infinite) sequence of pair of relations has a tree graph structure with two sequences outgoing from one point. At each point we call one branch ”left” and another branch as ”right” marking the corresponding pair by $L$ or $R$. We denote such pairs by $\left(\alpha_{i}^{(K_{i})}\right)$, where $\alpha_{i}^{(K_{i})}$ is an element of $C\left({n_{i}^{(K_{i})}}\right)$ involved in differential relations, and $\left(K_{i}\right)$ is a sequence of $i$ entries each is either $L$ or $R$ for $i\geq 1$.
Let $\chi\in C(n_{0})$, $\Phi\in C(n)$, for any $n_{0}$, $n\in\mathbb{Z}$. Due to the property (2.4) of the multiplication (2.2), the orthogonality condition (2.5) applied to $\Phi$ and $\chi$, i.e.,
@@
implies that there exists $\alpha_{1}^{(R)}\in C\left({n_{1}^{(R)}}\right)$, such that
@@
occurs. Acting by $\delta(n_{0}+1)$ on (2.7) we obtain
$0=\left(\delta(n)\Phi\right)\cdot\alpha^{(R)}_{1}+(-1)^{n_{0}+1}\Phi\cdot\delta\left(n^{(R)}_{1}\right)\alpha^{(R)}_{1}.$ (2.9)
On the other hand, (2.6) implies that there exists $\alpha_{1}^{(L)}\in C\left(n_{1}^{(L)}\right)$, such that
@@
where $r_{1}^{(l)}$ is the number of common parameters among $n_{1}^{(L)}$ and $(n_{0}+1)$ for $\alpha_{1}^{(L)}$ and $\delta(n_{0})\chi$.
Consequently applying the corresponding $\delta$ operators to (2.6), (2.7) and (2.10) we obtain the system of relations:
00 0 $=$ $\Phi\cdot\delta(n_{0})\chi,\qquad(1)$
@@
${\rm(1)}\;\Rightarrow\;\delta(n_{0})\chi$ $=$ $\Phi\cdot\alpha^{(R)}_{1}.$
00 0 $=$ $\delta(n)\Phi\cdot\alpha^{(R)}_{1}+(-1)^{n}\Phi\cdot\delta\left(n^{(R)}_{1}\right)\alpha^{(R)}_{1}\;\Rightarrow 0,$ (2.12)
[…]
$\Downarrow\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad(3^{\prime})$
@@
$\Downarrow\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad(3^{\prime\prime})$
@@
where we obtain (infinite) sequences (2.2) and (2.2) of pairs of relations for $\alpha_{i}^{(K_{i})}\in C\left({n_{i}^{(K_{i})}}\right)$, $i\geq 1$. Recall that we denote such relations as $\left(\alpha_{i}^{(K_{i})}\right)$. The corresponding indices $n_{i}^{(K_{i})}$ satisfy (in addition to (2.8) and (2.11)) relations for the sequence starting from $(4)$:

### v6

@@ Abstract.
Continual Lie algebras are infinite-dimensional generalizations of Lie algebras with discrete root system by considering continual root systems. In this paper we establish the general relation between chain complexes and continual Lie algebras. The natural orthogonality condition with respect to a product among elements of a chain complex $\mathcal{C}$ spaces brings about to $\mathcal{C}$ the structure of a graded algebra with differential relations. We prove the main result of this paper: a chain complex endowed with an appropriate Leibniz-property product of elements of its spaces and the Jacobi identity brings about the structure of a continual Lie algebra with the root space determined by parameters for the complex. That provides a new source of examples of continual Lie algebras. Finally, as an example, we consider the case of Čech-de Rham complex associated to a foliation of a smooth manifold. In a particular case of this chain complex, we derive explicitly the commutation relations for the corresponding continual Lie algebra.
@@ 1. Introduction
Continual Lie algebras introduced in [16] are generalizations of infinite-dimensional Lie algebras with systems of discrete roots [12]. Then the notion was generalized and developed in [17]-[20]. These algebras are formulated in such a way that the space of roots is defined by continual sets of vectors. In commutation relations, the generators depend on kernels which are functionals of continual roots. The Jacobi identity for continual Lie algebras result in non-trivial functional relations for kernels. Similar constructions appeared recently in the study of certain Hall algebras of coherent sheaves described in terms of certain continuum limits of Kac-Moody algebras [1].
Though the general theory of continual Lie algebras and their representations is missing, there exist many applications, especially in the theory of completely integrable and exactly solvable models [14]. Various applications can be also found in other fields of mathematics [4, 6, 20].
In the original papers [16]-[20], in order to find new classes of examples of continual Lie algebras various approaches were studied. In a series of papers [18, 19] the authors have generalized first known examples and have found new non-trivial and fundamental ones arising from various branches of mathematics, in particular, functional analysis [20], differential geometry, algebraic geometry, non-commutative geometry, and mathematical physics [4, 8]. Some of them came from physical applications, e.g., equations related to the Ricci flow and cosmology [4, 8]. Though they appear in various corners of modern mathematical research, at some stage, the arsenal of ideas of generation of new examples of continual Lie algebras was exhausted. An intriguing question is how to construct the general way which would allow to create continual Lie algebras. In many cases, this problem reduces to the question of finding new appropriate solutions for relations (1.1) (which is a very interesting problem by itself) following from the Jacobi identity.
The idea of this work is to use properties of the general chain complexes (with complex spaces depending on sets of parameters) endowed with a product in order to derive the structure of continual Lie algebras. The main result of this paper consists in the following
@@
For the set $I_{0}$ of all pairs of independent elements $\alpha_{i}^{(K_{i})}$, the orthogonality condition (LABEL:ortho), with existence of the left and right principal ideals with respect to subset of the complex, the generators $\left\{\chi,\;\delta(n_{0})\chi,\;\Phi,\;\delta(n)\Phi,\;\alpha_{i}^{(K_{i})},\;\delta(n(i))\alpha_{i}^{(K_{i})}\right\}$, $n(i)=n_{i}^{(K_{i})}$, the relations (2.2), form a graded differential algebra. If, depending on properties of the chain complex $\left(C(n,\Theta_{n}),\cdot\right)$, the kernels/mappings defining the commutation relations originating from the system (2.2) satisfy the Jacobi identity (1.1), then the bigraded differential algebra turns into a continual Lie algebra $\mathcal{G}(\Theta_{n(i)})$ with the root space depending on the set of parameters $\Theta_{n(i)}$, of spaces $C\left(n(i),\Theta_{n(i)}\right)$.
@@
The plan of the paper is the following. In the next subsection we recall the notion of a continual Lie algebra [16]. In Section ??, starting from the general (infinite) chain complex, we first define a product acting on elements of various spaces of a complex. We require the Leibniz rule to be fulfilled for (co)boundary operators with respect to this product. In addition to that we assume its skew-symmetry properties. Thus we endow the complex with the structure of a graded differential algebra $\mathcal{C}$.
@@
We then apply the natural condition of orthogonality with respect to a product among the chain complex spaces. For instance, in differential geometry, the orthogonality condition reduces to the integrability condition for differential forms. Further actions of (co)boundary operators, as well as consequences from the definition of the multiplication, generate a system of relations for elements of various chain complex spaces with extra compatibility conditions for indices. This determines the structure of a graded algebra with relations for $\mathcal{C}$. By choosing independent elements satisfying the above relations, we then show the main result of this paper: for a graded differential algebra associated to a chain complex, the independent generators, the system of relations, and the Jacobi identity define the structure of a continual Lie algebra with the root space provided by the spaces parameters of a chain complex.
@@
In Section ?? we consider the example of the Čech-de Rham cochain bicomplex [5] (c.f. subsection ??) associated to a smooth manifold foliation [7, 2, 3]. This bicomplex has a deep geometric meaning [9] and is defined for the spaces of differential forms and holonomy embeddings acting between sections of the transversal basis for a foliation. In subsection ?? we recall the notions of holonomy mappings, transversal sections, and transversal basis for a foliation. The space of differential forms has already a structure of bigraded differential algebra with respect to the natural product [5]. For the bicomplex under consideration, we define the second product, satisfying properties required for the construction of a system of relations. According to the exposition we give in Section ??, by involving the orthogonality conditions for an arbitrary pairs of the Čech-de Rham bicomplex, we apply the general scheme of construction of a bigraded differential algebra. Then we single out generators and commutation relations for the corresponding continual Lie algebra with the space of roots and kernels given by the sets of holonomy embeddings. In a particular geometric case associated to a codimension one foliation over a three-dimensional smooth manifold, we start from the integrability condition applied to elements of the same bicomplex space which leads to the appearance of a continual Lie algebra with generators, kernels, and commutation relations explicitly described.
@@ 1.1. Continual Lie algebras
In this subsection we recall the notion of a continual Lie algebra introduced in [17]. It was then studied in [18, 19]. Suppose $E$ is an associative algebra over $\mathbb{R}$ or $\mathbb{C}$, and $K_{0}$, $K_{\pm 1}$, $K_{0,0}:E\times E\to E$, are bilinear mappings. The local part of a continual Lie algebra can be defined as ${\widehat{\mathcal{G}}}=\mathcal{G}_{-1}\oplus\mathcal{G}_{0}\oplus\mathcal{G}_{+1}$, where $\mathcal{G}_{i}$, $i=0,\pm 1$. The subspaces $\mathcal{G}_{i}$ consist of the elements $\left\{X_{i}(\phi),\phi\in E\right\}$, $i=0,\pm 1$. The generators $X_{i}(\phi)$ are subject to commutation relations. For instance, in the principle grading, they have the form
@@
[…]
@@ 2.1. The graded differential algebra
In this section we formulate the construction which allows us to generate an algebra of graded relations starting from a chain complex. Consider a complex of spaces with elements depending on sets of parameters $\Theta_{i}$, $i\in\mathbb{Z}$, and given by
$\ldots\mathrel{{\mathop{\longrightarrow}\limits^{\delta(i-1)}}}C(i,\Theta_{i})\mathrel{{\mathop{\longrightarrow}\limits^{\delta(i)}}}C(i+1,\Theta_{i+1})\mathrel{{\mathop{\longrightarrow}\limits^{\delta(i+1)}}}\ldots,$ (2.1)
@@
so that for any elements $\Phi_{i}\in C(i,\Theta_{i})$ and $\Phi_{j}\in C(j,\Theta_{j})$, $\Phi_{i}\cdot\Phi_{j}=\Phi_{f(i,j)}$, where $f(i,j)$ and $g(i,j)$ are some functions of indices $i$ and $j$. Note that a function $f(i,j)$ defined the index for the product (LABEL:initialproduct) resulting space. The function $g(i,j)$ defines the index for the resulting set of parameters.
Let us describe the functions $f(i,j)$ and $g(i,j)$ that we use in this paper explicitly. For the elements $\Phi_{k}\in C(k,\Theta_{k})$, and $\Phi_{l}\in C(l,\Theta_{l})$, let $r$ be the number of common parameters among $k$- and $l$-sets of parameters for $\Phi_{k}$ and $\Phi_{l}$. We denote by $\deg\left(\Phi_{k}\right)=k$ the degree of an element $\Phi_{k}\in C(k,\Theta_{k})$. In addition to the properties of a product $\cdot$ mentioned above, we remove one set of $r$ common parameters from the product $\Phi_{k}\cdot\Phi_{l}$, thus, the total degree of the product $\Phi_{k}\cdot\Phi_{l}$ is $\deg(\Phi_{k}\cdot\Phi_{l})=k+l-r$. Let us assume that for all $i\in\mathbb{Z}$, the Leibniz rule formula for the operator $\delta(i)$ takes place, i.e.,
$\delta(k+l)\left(\Phi_{k}\cdot\Phi_{l}\right)=\left(\delta(k)\Phi_{k}\right)\cdot\Phi_{l}+(-1)^{\deg\left(\Phi_{k}\cdot\Phi_{l}\right)}\Phi_{k}\cdot\delta(l)\Phi_{l}.$ (2.3)
Note that in each part of the formula above, the total degree is k+lr+1k+l-r+1. We assume also that the chain complex above has some properties of an algebra with respect to the product (LABEL:initialproduct). Let us assume, in particular, that the product (LABEL:initialproduct) satisfies the condition
$\Phi_{k}\cdot\Phi_{l}=-(-1)^{\deg\left(\Phi_{k}\cdot\Phi_{l}\right)}\Phi_{l}\cdot\Phi_{k}.$ (2.4)
for all $k$, $l\in\mathbb{Z}$, $\Phi_{k}\in C(k,\Theta_{k})$ and $\Phi_{l}\in C(l,\Theta_{l})$. It is useful to introduce the following notation that we use in the text below: $\gamma_{n,m}=(-1)^{nm}$. Altogether, the complex (LABEL:complex) form a graded differential algebra $\mathcal{C}$. In what follows, we skip the sets of parameters $\Theta_{i}$ in the notations for the spaces of complexes, i.e., we set $C(i)=C(i,\Theta_{i})$. Though one has to keep in mind that a product of elements of two spaces of complex (LABEL:complex) depends on the resulting set of parameters $\Theta_{g(i,j)}$.
@@ 2.2. The algebra of graded differential relations
As stated above, we assume that the product (LABEL:initialproduct) is defined in such a way that common parameters of the sets $\Theta_{i}$ and $\Theta_{j}$ are present in the resulting set $\Theta_{g(i,j)}$ only once. For some $i$ and $j\in\mathbb{Z}$, let us introduce the orthogonality condition for a pair $C(i,\Theta_{i})$ and $C(j,\Theta_{j})$, with respect to the product (LABEL:initialproduct). In particular, let us require that for a pair $C(i,\Theta_{i})$, $C(j,\Theta_{j})$, there exist subspaces $C^{\prime}(i)\subset C(i,\Theta_{i})$ and $C^{\prime}(j)\subset C(j,\Theta_{j})$, such that, for any $\Phi_{i}\in C^{\prime}(i)$ and $\Phi_{j}\in C^{\prime}(j)$,
@@
namely, $\Phi_{i}$ is supposed to be orthogonal to $\delta(j)\Phi_{j}$ with respect to (LABEL:initialproduct).
By applying further differentials to (LABEL:ortho) (and further consequences of such action), and using properties of a particular function $f(i,j)$ we obtain relations among elements of spaces $C(i,\Theta_{i})$, $i\in\mathbb{Z}$. In particular, taking into account that both sides of such relations belong to the same space, we obtain limitations (depending on the function $f$) on indices. In differential geometry, the orthogonality condition (LABEL:ortho) provides the definition of integrability conditions for differential forms, and leads to the Frobenius theorem (see, e.g., [10]).
Let us explain the notations we will use on few next pages. In (2.2) we obtain (infinite) sequences of pairs of relations of the form (LABEL:trojka)-(LABEL:trojka1). The (infinite) sequence of pair of relations has a tree graph structure with two sequences outgoing from one point. At each point we call one branch ”left” and another branch as ”right” marking the corresponding pair by $L$ or $R$. We denote such pairs by $\left(\alpha_{i}^{(K_{i})}\right)$, where $\alpha_{i}^{(K_{i})}$ is an element of $C\left({n_{i}^{(K_{i})}}\right)$ involved in relations, and $\left(K_{i}\right)$ is a sequence of $i$ entries each is either $L$ or $R$ for $i\geq 1$.
Let $\chi\in C(n_{0})$, $\Phi\in C(n)$, for any $_{0}, nn\in\mathbb{Z}$. Assume that for each $\Phi$ the annihilator $\mathrm{Ann}_{R}(\Phi)=\left\{\phi\in C(i,\Theta_{i})\mid\Phi\cdot\phi=0\right\}$ is a principal right ideal, i.e., there exists an element $\alpha^{(R)}_{1}$ such that $\mathrm{Ann}_{R}(\Phi)=\left\{\Phi\cdot C(j,\Theta_{j})\right\}$. Then the orthogonality condition(LABEL:ortho) applied to $\Phi$ and $\chi$ implies, i.e.,
@@
$\delta(n_{0})\chi\in\mathrm{Ann}_{R}(\Phi)=\left\{\Phi\cdot C(k,\Theta_{k})\right\}$, implies that there exists $\alpha^{(R)}_{1}\in C\left({n_{1}^{(R)}}\right)$ such that
@@
occurs. Acting by $\delta(n_{0}+1)$ on (LABEL:trojka) we obtain
$0=\left(\delta(n)\Phi\right)\cdot\alpha^{(R)}_{1}+(-1)^{(n_{0}+1)n_{1}^{(R)}}\Phi\cdot\delta\left(n^{(R)}_{1}\right)\alpha^{(R)}_{1}$. (2.9)
Similarly, if the left annihilator $\mathrm{Ann}_{L}(\delta(n_{0})\chi)=\left\{\psi\mid\psi\cdot\delta(n_{0})\chi=0\right\}$ is principal, then (LABEL:introrth0) implies that there exists $\alpha_{1}^{(L)}\in C\left(n_{1}^{(L)}\right)$, such that
@@
where $r_{1}^{(l)}$ is the number of common parameters among $n_{1}^{(L)}$ and $(n_{0}+1)$ for $\alpha_{1}^{(L)}$ and $\delta(n_{0})\chi$. Consequently applying the corresponding $\delta$ operators to (LABEL:introrth0), (LABEL:trojka) and (2.10) we obtain the system of relations:
$0$ $=$ $\Phi\cdot\delta(n_{0})\chi,\qquad(1)$
@@
${\delta.\rm(1)}\;\Rightarrow\;0$ $=$ $\left(\delta(n)\Phi\right)\cdot\delta(n_{0})\chi,\qquad(3)$
${\rm(1)}\;\Rightarrow\;\delta(n_{0})\chi$ $=$ $\Phi\cdot\alpha^{(R)}_{1}$.
$0$ $=$ $\delta(n^{(R)}_{1})\Phi\cdot\alpha^{(R)}_{1}+(-1)^{nn^{(R)}_{1}}\Phi\cdot\delta\left(n^{(R)}_{1}\right)\alpha^{(R)}_{1}\;\Rightarrow 0,$ (2.12)
${\rm(4):}\;\;0$ $=$ $\left(\delta(n)\Phi\right)\cdot\delta\left(n_{2}^{(RR)}\right)\alpha^{(RR)}_{2}\Rightarrow\left(\alpha^{(RRR)}_{3}\right)\Rightarrow$... (2.13)
@@
${\rm(5):}\;\;0$ $=$ $\delta\alpha^{(LL)}_{2}\cdot\delta(n_{0})\chi\Rightarrow\left(\alpha^{(LLR)}_{3}\right)\Rightarrow$... (2.14)
@@
where we obtain (infinite) sequences (2.2) and (2.2) of pairs of relations for $\alpha_{i}^{(K_{i})}\in C\left({n_{i}^{(K_{i})}}\right)$, $i\geq 1$. Recall that we denote such relations as $\left(\alpha_{i}^{(K_{i})}\right)$. The corresponding indices $n_{i}^{(K_{i})}$ satisfy (in addition to (LABEL:condi1) and (LABEL:condi2)) relations for the sequence starting from $(4)$: